# From Similarity to Superiority: Channel Clustering for Time Series Forecasting

### v1

@@ 5.6 Ablation Study on Cluster Number
Figure 4 shows an ablation study on cluster ratios, which is defined as the ratio of the number of clusters to the number of channels. $0.0$ means all channels are in a single cluster. We observe that the MSE loss slightly decreases and then increases as the cluster ratio increases, especially for DLinear, PatchTST, and TimesNet. Time series models with CCM achieve the best performance when the cluster ratio is in the range of $[0.2,0.6]$. It is worth noticing that DLinear and PatchTST, two CI models among four base models, benefit consistently from channel clustering with any number of clusters. More results are provided in Appendix D.
@@ 6 Conclusion
This work introduces a novel Channel Clustering Module (CCM) to address the challenge of effective channel management in time series forecasting. CCM strikes a balance between individual channel treatment and capturing cross-channel dependencies by clustering channels based on their intrinsic similarity. Extensive experiments demonstrate the efficacy of CCM in multiple benchmarks. It would be valuable to investigate the applicability of CCM in other domains beyond time series forecasting in future works.
Impact Statement. This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Channel Similarity
Essentially, the similarity between two time series $X_{i}$ and $X_{j}$ is defined as $\textsc{SIM}(X_{i},X_{j})=\text{exp}(\frac{-d(X_{i},X_{j})}{2\sigma^{2}})$, where $d(\cdot,\cdot)$ can be any distance metric (Chen & Ng, 2004; Ding et al., 2008), such as Euclidean Distance ($L_{p}$), Editing Distance (ED) and Dynamic Time Warping (DTW) (Gold & Sharir, 2018). One may also use other similarity definitions, such as Longest Common Subsequence (LCSS) and Cross-correlation (CCor).
@@ Appendix A Channel Similarity
Secondly, it’s worth noting that while there are various similarity computation approaches, studies have demonstrated a strong correlation between Euclidean distance and other distance metrics (Kljun & Tersˇek, 2020). This high correlation suggests that, despite different mathematical formulations, these metrics often yield similar results when assessing the similarity between time series. This empirical evidence supports the choice of Euclidean distance as a reasonable approximation of similarity for practical purposes. In our implementation, we select $\sigma=5$ in Eq. 1 to compute the similarities based on Euclidean distance.
@@ Appendix B Multivariate and Univariate Adaptation
  • Training on a multivariate dataset and inferring on either univariate or multivariate samples: For Channel-Dependent models, test samples should have the same number of channels as the training samples, as seen in sub-datasets within ETT collections (Zhou et al., 2021). In contrast, for Channel-Independent models, zero-shot forecasting can be performed on either univariate or multivariate samples, even when they have different numbers of channels.
@@ C.1 Datasets
Public Datasets. We utilize nine commonly used datasets for long-term forecasting evaluation. Firstly, ETT collection (Zhou et al., 2021), which documents the oil temperature and load features of electricity transformers over the period spanning July 2016 to July 2018. This dataset is further subdivided into four sub-datasets, ETThs and ETTms, with hourly and 15-minute sampling frequencies, respectively. s can be $1$ or $2$, indicating two different regions. Electricity dataset (Trindade, 2015) encompasses electricity consumption data from 321 clients from July 2016 to July 2019. Exchange (Lai et al., 2017) compiles daily exchange rate information from 1990 to 2016. Traffic dataset contains hourly traffic load data from 862 sensors in San Francisco areas from 2015 to 2016. Weather dataset offers a valuable resource with 21 distinct weather indicators, including air temperature and humidity, collected every 10 minutes throughout the year 2021. ILI documents the weekly ratio of influenza-like illness patients relative to the total number of patients, spanning from 2002 to 2021. Dataset statistics can be found in Table 10.
@@ C.2 Metrics
Following standard evaluation protocols (Wu et al., 2022), we utilize the Mean Absolute Error (MAE) and Mean Square Error(MSE) for long-term and stock price forecasting. The Symmetric Mean Absolute Percentage Error (SMAPE), Mean Absolute Scaled Error (MASE), and Overall Weighted Average (OWA) are used as metrics for M4 dataset (Makridakis et al., 2018; Oreshkin et al., 2019). The formulations are shown in Eq. 5. Let $\mathbf{y}_{t}$ and $\widehat{\mathbf{y}}_{t}$ denote the ground-truth and the forecast at the $t$-th timestep, respectively. $H$ is the forecasting horizon. In M4 dataset, MASE is a standard scale-free metric, where $s$ is the periodicity of the data (e.g., 12 for monthly recorded sub-dataset) (Makridakis et al., 2018). MASE measures the scaled error w.r.t. the naïve predictor that simply copies the historical records of $s$ periods in the past. Instead, SMAPE scales the error by the average between the forecast and ground truth. Particularly, OWA is an M4-specific metric (Team et al., 2018) that assigns different weights to each metric.
@@
OWA=12[ $\text{OWA}=\frac{1}{2}\left[\frac{\text{ SMAPE }}{\text{ SMAPE }_{\text{Naïve2 }}}+\frac{\text{ MASE }}{\text{MASE}_{\text{Naïve2 }}}\right]$
@@ C.3 Experiment Details
To verify the superiority of CCM in enhancing the performance of mainstream time series models, we select four popular and state-of-the-art models, including linear models such as TSMixer (Chen et al., 2023), DLinear (Zeng et al., 2022), transformer-based model PatchTST (Nie et al., 2022) and convolution-based model TimesNet (Wu et al., 2022). We build time series models using their official codes and optimal model configuration [^1] https://github.com/yuqinie98/PatchTST [^2] https://github.com/cure-lab/LTSF-Linear [^3] https://github.com/google-research/google-research/tree/master/tsmixer [^4] https://github.com/thuml/TimesNet.
In the data preprocessing stage, we apply reversible instance normalization (Kim et al., 2021) to ensure zero mean and unit standard deviation, avoiding the time series distribution shift. Forecasting loss is MSE for long-term forecasting datasets and the stock dataset. Instead, we use SMAPE loss for M4 dataset. We select Adam (Kingma & Ba, 2014) with the default hyperparameter configuration for $(\beta_{1},\beta_{2})$ as (0.9, 0.999). An early-stopping strategy is used to mitigate overfitting. The experiments are conducted on a single NVIDIA RTX A6000 48GB GPU, with PyTorch (Paszke et al., 2019) framework. The code is available at the following anonymous link: https://anonymous.4open.science/r/TimeSeriesCCM-ECC2. We use the official codes and follow the best model configuration to implement the base models. Then we apply CCM to the base models, keeping the hyperparameters unchanged for model backbones. Experiment configurations, including the number of MLP layers in the cluster assigner, the layer number in the temporal module, hidden dimension, the best cluster number, and regularization parameter $\beta$ on nine real-world datasets are shown in Table 11.
@@ C.4 Comparison between CCM and Other Approach
Predict Residuals with Regularization. Prior work (Han et al., 2023) demonstrates that the main drawback of CD models is their inclination to generate sharp and non-robust forecasts, deviating from the actual trend. Thereby, Predict Residuals with Regularization (PRReg for simplicity), a specifically designed regularization objective, is proposed to improve the robustness of CD methods as follows.
@@
where the superscript $j$ indicates the sample index. $\mathcal{L}_{F}$ is MSE loss. $A^{(j)}=X^{j}_{L,:}$ represents the last values of each channel in $X^{(j)}$. Therefore, the objective encourages the model to generate predictions that are close to the nearby historical data and keep the forecasts smooth and robust. The regularization term $\Omega$ ($L_{2}$ norm in practice) further restricts the complexity of the model and ensures smoothness in the predictions. It was demonstrated that PRReg can achieve better performance than original CD and CI strategies (Han et al., 2023). We conduct extensive experiments on long-term forecasting benchmarks to compare PRReg and CCM. The full results are shown in Table 12. The numbers in parentheses next to the method represent the forecasting horizon. We set the length of the look-back window to 36 for ILI and 96 for other datasets for consistency with previous works (Han et al., 2023). The base models are linear model (Zeng et al., 2022) and basic transformer (Vaswani et al., 2017). The values in the PRReg column are the best results with any $\lambda$, collected from (Han et al., 2023). We observe from Table 12 that CCM successfully improves forecasting performance on original CI/CD strategies (or reached comparable results) in 13 out of 16 settings, compared with PRReg method.
(a) Model Size Analysis on ETTh1 dataset
(b) Runtime Efficiency Analysis on ETTh1 dataset
C.5 Complexity and Efficiency Analysis
@@ C.6 Error Bar
Table 15: Standard deviation of Table  7 on stock dataset
@@ D.1 Influence of Cluster Ratio
The number of clusters is an important hyperparameter in the CCM method. To verify the effectiveness of our design, we conduct an ablation study for all base models on four long-term forecasting datasets. The full results are shown in Table 16. We tune different cluster ratios, defined as the ratio of the number of clusters to the number of channels. Original means the original base model without any channel clustering mechanism. 0.0 indicates grouping all channels into the same cluster. We make the following observations. (1) For most cases, the channel clustering module (CCM) with any number of clusters greater than 1 consistently improves the forecasting performance upon base models. (2) For Channel-Independent base models, such as DLinear and PatchTST, grouping all channels into one cluster results in a channel fusion at the last layer, leading to a degradation in forecasting performance than the original CI models. (3) For most cases, the cluster ratio in the range of $[0.2,0.6]$ typically results in the best performance.

### v2

@@ 5.6 Ablation Studies
Figure 4 shows an ablation study on cluster ratios, which is defined as the ratio of the number of clusters to the number of channels. $0.0$ means all channels are in a single cluster. We observe that the MSE loss slightly decreases and then increases as the cluster ratio increases, especially for DLinear, PatchTST, and TimesNet. Time series models with CCM achieve the best performance when the cluster ratio is in the range of $[0.2,0.6]$. It is worth noticing that DLinear and PatchTST, two CI models among four base models, benefit consistently from channel clustering with any number of clusters. Additional ablation studies on the look-back window length and clustering step are provided in Appendix D.
5.7 Efficiency Analysis
(a) Model Size Analysis on ETTh1 dataset
(b) Runtime Efficiency Analysis on ETTh1 dataset
@@ 6 Conclusion
This work introduces a novel Channel Clustering Module (CCM) to address the challenge of effective channel management in time series forecasting. CCM strikes a balance between individual channel treatment and capturing cross-channel dependencies by clustering channels based on their intrinsic similarity. Extensive experiments demonstrate the efficacy of CCM in multiple benchmarks, including long-term, short-term, and zero-shot forecasting scenarios. Refinement of the CCM clustering and domain-specific similarity measurement could potentially improve the model performance further. Moreover, it would be valuable to investigate the applicability of CCM in other domains beyond time series forecasting in future works.
Acknowledgments and Disclosure of Funding
This research was supported in part by the National Science Foundation (NSF) CNS Division Of Computer and Network Systems (2431504) and AWS Research Awards. We would like to thank the anonymous reviewers for their constructive feedback. Their contributions have been invaluable in facilitating our work.
Appendix A Definitions
A.1 Channel Similarity
Essentially, the similarity between two time series $X_{i}$ and $X_{j}$ is defined as $\textsc{SIM}(X_{i},X_{j})=\text{exp}(\frac{-d(X_{i},X_{j})}{2\sigma^{2}})$, where $d(\cdot,\cdot)$ can be any distance metric [74, 75], such as Euclidean Distance ($L_{p}$), Editing Distance (ED) and Dynamic Time Warping (DTW) [76]. One may also use other similarity definitions, such as Longest Common Subsequence (LCSS) and Cross-correlation (CCor).
@@ A.1 Channel Similarity
Secondly, it’s worth noting that while there are various similarity computation approaches, studies have demonstrated a strong correlation between Euclidean distance and other distance metrics [77]. This high correlation suggests that, despite different mathematical formulations, these metrics often yield similar results when assessing the similarity between time series. This empirical evidence supports the choice of Euclidean distance as a reasonable approximation of similarity for practical purposes. In our implementation, we select $\sigma=5$ in Eq. 1 to compute the similarities based on Euclidean distance.
A.2 Channel Dependent and Channel Independent Strategy
The definitions of Channel Dependent (CD) and Channel Independent (CI) settings are pivotal to this work. The fundamental difference lies in whether a model captures cross-channel information. There are slightly varied interpretations of Channel Independent (CI) in previous works and we summarize as follows.
  1. 1.
    In some works [15, 54], CI is broadly defined as forecasting each channel independently, where cross-channel dependencies are completely ignored. For linear models [78, 16], CI is specifically defined as $n$ individual linear layers for $n$ channels in previous works. Each linear layer is dedicated to modeling a univariate sequence, with the possibility of differing linear weights across channels.
  2. 2.
    In previous work [17], CI also means all channels being modeled independently yet under a unified model.
All the above works acknowledge that CI strategies often outperform CD approaches, though this comparison is not the focal point of our work. It’s also recognized that the specific CI strategy employed in DLinear and PatchTST contributes significantly to their performance. The CI setting in [17] represents a specific instance within the broader CI setting in other works [15, 54]. To avoid ambiguity, we use $f^{(i)}$ to represent the model for the $i$-th channel specifically, aligning with previous definitions without conflict.
@@ Appendix B Multivariate and Univariate Adaptation
  • Training on a multivariate dataset and inferring on either univariate or multivariate samples: For Channel-Dependent models, test samples should have the same number of channels as the training samples, as seen in sub-datasets within ETT collections [11]. In contrast, for Channel-Independent models, zero-shot forecasting can be performed on either univariate or multivariate samples, even when they have different numbers of channels.
@@ C.1 Datasets
Public Datasets. We utilize nine commonly used datasets for long-term forecasting evaluation. Firstly, ETT collection [11], which documents the oil temperature and load features of electricity transformers over the period spanning July 2016 to July 2018. This dataset is further subdivided into four sub-datasets, ETThs and ETTms, with hourly and 15-minute sampling frequencies, respectively. s can be $1$ or $2$, indicating two different regions. Electricity dataset [79] encompasses electricity consumption data from 321 clients from July 2016 to July 2019. Exchange [80] compiles daily exchange rate information from 1990 to 2016. Traffic dataset contains hourly traffic load data from 862 sensors in San Francisco areas from 2015 to 2016. Weather dataset offers a valuable resource with 21 distinct weather indicators, including air temperature and humidity, collected every 10 minutes throughout the year 2021. ILI documents the weekly ratio of influenza-like illness patients relative to the total number of patients, spanning from 2002 to 2021. Dataset statistics can be found in Table 9.
@@ C.2 Metrics
Following standard evaluation protocols [13], we utilize the Mean Absolute Error (MAE) and Mean Square Error(MSE) for long-term and stock price forecasting. The Symmetric Mean Absolute Percentage Error (SMAPE), Mean Absolute Scaled Error (MASE), and Overall Weighted Average (OWA) are used as metrics for M4 dataset [72, 42]. The formulations are shown in Eq. 5. Let $\mathbf{y}_{t}$ and $\widehat{\mathbf{y}}_{t}$ denote the ground-truth and the forecast at the $t$-th timestep, respectively. $H$ is the forecasting horizon. In M4 dataset, MASE is a standard scale-free metric, where $s$ is the periodicity of the data (e.g., 12 for monthly recorded sub-dataset) [72]. MASE measures the scaled error w.r.t. the naïve predictor that simply copies the historical records of $s$ periods in the past. Instead, SMAPE scales the error by the average between the forecast and ground truth. Particularly, OWA is an M4-specific metric [81] that assigns different weights to each metric.
@@
OWA=12[ $\text{OWA}=\frac{1}{2}\left[\frac{\text{ SMAPE }}{\text{ SMAPE }_{\text{Na\"{\i}ve2 }}}+\frac{\text{ MASE }}{\text{MASE}_{\text{Na\"{\i}ve2 }}}\right]$
@@ C.3 Experiment Details
To verify the superiority of CCM in enhancing the performance of mainstream time series models, we select four popular and state-of-the-art models, including linear models such as TSMixer [7], DLinear [8], transformer-based model PatchTST [21] and convolution-based model TimesNet [13]. We build time series models using their official codes and optimal model configuration [^2] https://github.com/yuqinie98/PatchTST [^3] https://github.com/cure-lab/LTSF-Linear [^4] https://github.com/google-research/google-research/tree/master/tsmixer [^5] https://github.com/thuml/TimesNet.
In the data preprocessing stage, we apply reversible instance normalization [64] to ensure zero mean and unit standard deviation, avoiding the time series distribution shift. Forecasting loss is MSE for long-term forecasting datasets and the stock dataset. Instead, we use SMAPE loss for M4 dataset. We select Adam [82] with the default hyperparameter configuration for $(\beta_{1},\beta_{2})$ as (0.9, 0.999). An early-stopping strategy is used to mitigate overfitting. The experiments are conducted on a single NVIDIA RTX A6000 48GB GPU, with PyTorch [83] framework. We use the official codes and follow the best model configuration to implement the base models. Then we apply CCM to the base models, keeping the hyperparameters unchanged for model backbones. Experiment configurations, including the number of MLP layers in the cluster assigner, the layer number in the temporal module, hidden dimension, the best cluster number, and regularization parameter $\beta$ on nine real-world datasets are shown in Table 10.
@@ C.4 Comparison between CCM and Other Approach
Predict Residuals with Regularization. Prior work [17] demonstrates that the main drawback of CD models is their inclination to generate sharp and non-robust forecasts, deviating from the actual trend. Thereby, Predict Residuals with Regularization (PRReg for simplicity), a specifically designed regularization objective, is proposed to improve the robustness of CD methods as follows.
@@
where the superscript $j$ indicates the sample index. $\mathcal{L}_{F}$ is MSE loss. $A^{(j)}=X^{j}_{L,:}$ represents the last values of each channel in $X^{(j)}$. Therefore, the objective encourages the model to generate predictions that are close to the nearby historical data and keep the forecasts smooth and robust. The regularization term $\Omega$ ($L_{2}$ norm in practice) further restricts the complexity of the model and ensures smoothness in the predictions. It was demonstrated that PRReg can achieve better performance than original CD and CI strategies [17]. We conduct extensive experiments on long-term forecasting benchmarks to compare PRReg and CCM. The full results are shown in Table 11. The numbers in parentheses next to the method represent the forecasting horizon. We set the length of the look-back window to 36 for ILI and 96 for other datasets for consistency with previous works [17]. The base models are linear model [8] and basic transformer [41]. The values in the PRReg column are the best results with any $\lambda$, collected from [17]. We observe from Table 11 that CCM successfully improves forecasting performance on original CI/CD strategies (or reached comparable results) in 13 out of 16 settings, compared with PRReg method.
C.5 Results Analysis
We report the degree of multivariate correlation across multiple channels (measured by the average Pearson correlation coefficient) in Table 12 and Table 13. $r$ denotes the degree of multivariate correlation. Then the Pearson correlation coefficient between $r$ and the performance improvement rate is $0.258$ in long-term forecasting, indicating a weak correlation. It demonstrates that CCM tends to achieve a greater boost on datasets that are intrinsically correlated within channels. Compared with datasets used in long-term benchmarks, M4 demonstrates more significant correlations between time series. Therefore, capturing the correlation within the dataset in short-term cases potentially leads to greater improvement in the forecasting performance than in long-term cases.
Table 12: Multivariate intrinsic similarity for long-term forecasting datasets
Dataset ETTh1 ETTm1 ETTh2 ETTm2 Exchange ILI Weather Electricity Traffic
Correlation $r$ 0.1876 0.1717 0.3224 0.328 0.3198 0.508 0.1169 0.5311 0.6325
Table 13: Intrinsic similarity for short-term forecasting datasets
Dataset M4 Monthly M4 Daily M4 Yearly M4 Hourly M4 Quarterly M4 Weekly
Correlation $r$ 0.62 0.646 0.712 0.55 0.671 0.653
@@ C.6 Error Bar
Table 16: Standard deviation of Table  6 on stock dataset
@@ D.1 Influence of Cluster Ratio
The number of clusters is an important hyperparameter in the CCM method. To verify the effectiveness of our design, we conduct an ablation study for all base models on four long-term forecasting datasets. The full results are shown in Table 17. We tune different cluster ratios, defined as the ratio of the number of clusters to the number of channels. Original means the original base model without any channel clustering mechanism. 0.0 indicates grouping all channels into the same cluster. We make the following observations. (1) For most cases, the channel clustering module (CCM) with any number of clusters greater than 1 consistently improves the forecasting performance upon base models. (2) For Channel-Independent base models, such as DLinear and PatchTST, grouping all channels into one cluster results in a channel fusion at the last layer, leading to a degradation in forecasting performance compared to the original CI models. (3) For most cases, the cluster ratio in the range of $[0.2,0.6]$ typically results in the best performance.
@@ D.2 Influence of Look-back Window Length
D.3 Influence of Different Clustering Steps
We conducted an ablation study on different clustering steps to investigate its effect on downstream performance, reported in Table 19. The ablation study follows the setting in Table 5. Random means we randomly assign each channel to a cluster, leading to worse clustering quality. K-Means means using the k-means algorithm to replace our clustering step, leading to suboptimal prototype embedding. The proposed CCM is essentially an advanced variant of K-Means with learnable prototype embedding and cross-attention mechanism. Results show that both clustering quality and prototype embedding will affect the downstream performance. Instead, CCM generates high-quality channel clustering results, compared with random assignment and K-Means clustering.
Table 19: Ablation on different clustering steps on ETTh1 and ETTm1 based on Linear and Transformer architecture.
CD CI Random K-Means CCM
ETTh1 Linear 0.402 0.345 0.389 0.357 0.342
Transformer 0.861 0.655 0.746 0.542 0.518
ETTm1 Linear 0.404 0.354 0.371 0.326 0.310
Transformer 0.458 0.379 0.428 0.311 0.300
@@ Appendix F Discussion
Social Impact. The Channel Clustering Module (CCM) presented in this paper holds significant potential for positive social impact. By improving the accuracy and efficiency of forecasting, CCM can benefit a wide range of applications critical to society. For instance, in healthcare, CCM could enhance the prediction of patient outcomes, leading to better treatment planning and resource allocation. Additionally, in financial markets, CCM could aid in predicting market trends, supporting informed decision-making and potentially reducing economic risks for individuals and organizations. Overall, the development and refinement of CCM could potentially enhance the quality of life and promote societal well-being.